# From Prompt Injections to Protocol Exploits: Threats in LLM-Powered AI Agents Workflows

### v1

@@ III Input Manipulation Attacks
This section groups techniques that subvert LLM behavior by tampering with the inputs, whether text, code, or multimodal data, before they reach the core model. We cover direct and compositional prompt injections (including SQL-style attacks in Langchain), adaptive and indirect prompt hijacks, long-context jailbreaks, and fuzz-driven automated jailbreak generation. Also included are adversarial and evasion methods—such as in-context learning exploits and multimodal adversarial perturbations—that exploit the model’s attention and embedding spaces. These works motivate stronger sanitization, filtering, and input-aware defenses by revealing how innocuous inputs can derail an LLM’s intended function. Tab. III presents a summary of research works on input manipulation attacks. Fig. 3 organizes input manipulation attacks into three broad categories: Prompt-Based, Jailbreaking, and Adversarial & Evasion.
III-A Prompt-Based Attacks
@@
III-A1 Direct Prompt Injection
@@
Perez et al. [33] investigate vulnerabilities in transformer-based large language models (LLMs), specifically GPT-3, that arise from malicious user interactions. Despite the widespread adoption of such models in customer-facing applications, there is a lack of research on their susceptibility to adversarial manipulation. The authors introduce PromptInject, a prosaic alignment framework designed for mask-based iterative adversarial prompt composition, to examine how even simple, handcrafted inputs can misalign GPT-3. The study focuses on two primary attack types: goal hijacking, where the attacker manipulates the model to perform unintended actions, and prompt leaking, where sensitive information is inadvertently exposed through the model’s responses. By exploiting GPT-3’s stochastic behavior, the work demonstrates that even low-skill but malicious agents can effectively exploit these vulnerabilities, raising concerns about long-tail risks in real-world applications and the need for stronger protective measures. Fig. 4 illustrates the adversarial inputs designed to manipulate the instructions provided to a language model, which is presented in [33]. Both depicted attacks aim to subvert the initial intent of the prompt. In the objective hijacking scenario, the attacker intends for the model to output a predefined malicious phrase, thus overriding the original instructions. Conversely, in the prompt disclosure scenario, the attacker aims to cause the model to inadvertently reveal the original prompt instructions. The Original Application Instructions contain placeholders user_provided{user\_provided}, representing regular user inputs intended for correction. Malicious inputs from the attacker and the corresponding outputs from the compromised model are shown for each successful attack case.
III-A2 Prompt-to-SQL (P2SQL) injection attack
@@
Pedro et al. [34] address the security vulnerabilities in web applications utilizing Large Language Models (LLMs), specifically focusing on SQL injection attacks facilitated by unsanitized user prompts. In systems where LLMs are integrated with middleware such as Langchain, user input is converted into SQL queries that guide the model’s responses. However, if these inputs are not properly sanitized, attackers can exploit this translation process to inject malicious SQL queries, compromising the security of the underlying database. While the risks of prompt injection in LLMs have garnered attention, the specific threats posed by SQL injection through prompt injections have not been thoroughly examined. The authors present a comprehensive study on prompt-to-SQL (P2SQL) injections in Langchain-based web applications. Through detailed case studies, the paper characterizes various forms of P2SQL attacks and evaluates the impact of these vulnerabilities on application security. Additionally, the paper assesses seven state-of-the-art LLMs, revealing the widespread susceptibility to P2SQL attacks. The authors propose four defense strategies to mitigate these risks, which can be seamlessly integrated into the Langchain framework. These defenses are validated through experimental evaluations, highlighting their efficacy in protecting against P2SQL injection attacks in real-world applications.
III-A3 Indirect & Compositional Prompt Injection
@@
By manipulating images and sounds, Bagdasaryan et al. [35] explore the potential for indirect prompt and instruction injection in multimodal large language models (LLMs). In this approach, an attacker creates an adversarial perturbation that aligns with a specific prompt or instruction. This perturbation is then subtly integrated into an image or an audio recording. When the user interacts with the model using the unaltered image or audio, the embedded perturbation influences the model’s output, effectively steering the conversation or responses in line with the attacker’s preferences. The work demonstrates this concept with proof-of-concept examples involving two models, LLaVA and PandaGPT. This form of attack highlights the vulnerability of multimodalmodels to adversarial inputs, emphasizing the need for enhanced defenses to detect and mitigate such manipulations in real-world applications.
III-A4 Adaptive indirect prompt injection attack
@@
Zhan et al. [16] examine the security risks posed by indirect prompt injection (IPI) attacks in Large Language Model (LLM) agents that utilize external tools to interact with their environments. While various defenses have been proposed to mitigate IPI attacks, their effectiveness has not been adequately tested against adaptive attacks. The authors evaluate eight existing defense mechanisms and demonstrate that all of them can be bypassed through adaptive attack strategies, resulting in an attack success rate exceeding 50% in each case. This highlights significant vulnerabilities in current defense approaches and emphasizes incorporating adaptive attack evaluation when designing security measures for LLM agents. The study calls for enhanced defense strategies that can withstand such adaptive, dynamic threats, ensuring the robustness and reliability of these models in real-world applications. Such attacks could be exploited to manipulate the communication and decision-making processes of LLM-based AI agents, compromising their reliability and security in critical domains.
III-A5 Toxic Agent Flow attack
@@
Toxic Agent Flow attack is a critical vulnerability in the widely-used GitHub MCP server, discovered by Invariant Labs  [^1] https://invariantlabs.ai/blog/mcp-github-vulnerability., that allows attackers to hijack a user’s agent through malicious GitHub issues, leading to the leakage of private repository data. As presented in Fig. 5, the vulnerability exploits a Toxic Agent Flow, a scenario where agents are manipulated into performing unintended actions, such as leaking sensitive data, through seemingly benign instructions. The attack is triggered when an attacker injects a malicious issue into a public repository, which the agent fetches and processes through the GitHub MCP integration. Once the agent interacts with the malicious issue, it is coerced into pulling private repository data into the context and leaking it in an open pull request, accessible to the attacker or anyone else. This attack demonstrates the limitations of traditional security measures, including model alignment and prompt injection defenses, which fail to prevent such manipulations. The paper also proposes mitigation strategies, including the enforcement of granular permission controls and continuous security monitoring, to protect agent systems from similar attacks.
@@ III-B Jailbreaking Attacks
The attack formulation [36] focuses on prompting the model to predict the next token in a sequence by using a probability distribution
@@
III-B1 Compositional Instruction Attack
@@
The recent study by Jiang et al. [17] on Compositional Instruction Attacks (CIA) presents a novel framework to better understand jailbreak vulnerabilities in large language models (LLMs), particularly in critical applications such as autonomous driving (AD). The researchers introduce two new methods for automatically generating tailored jailbreak prompts, which are designed to bypass LLM security mechanisms by embedding harmful instructions within seemingly harmless ones. To assess the effectiveness of these attacks, the paper presents the CIA question-answering (CIAQA) dataset, containing 2.7K multiple-choice questions based on 900 successful jailbreaks. The experimental results reveal that these attacks can achieve a high attack success rate (ASR) of 95% or higher on popular LLMs, such as GPT-4, GPT-3.5, and Llama2-70b-chat. The study also identifies key reasons for LLM defenses failing against such attacks, primarily due to the inability of LLMs to properly discern underlying harmful intents and prioritize tasks effectively. To counter these threats, the authors propose an intent-based defense paradigm (IBD), which leverages LLMs’ ability to identify harmful intents and significantly mitigates the attack success rate by over 74%. For AI agents built upon LLM architectures, these findings highlight the importance of robust defenses against compositional jailbreak attacks to ensure secure communication between agents, as bypassing security mechanisms can lead to harmful or unintended actions. As LLMs are integrated into autonomous agent communication protocols, similar attack vectors could potentially be exploited, emphasizing the need for stronger safeguards to protect against malicious input manipulations.
III-B2 Jailbreak via In-Context Demonstrations
@@
Wei et al. [36] explore the use of In-Context Learning (ICL) as a mechanism to influence the alignment of Large Language Models (LLMs), focusing on mitigating safety concerns related to harmful content generation. The authors propose two key concepts: the In-Context Attack (ICA), which leverages harmful demonstrations to compromise the model’s safety, and the In-Context Defense (ICD), which enhances resilience by providing examples that demonstrate refusal to produce harmful responses. The paper presents a theoretical framework that illustrates how a small set of carefully chosen in-context demonstrations can significantly alter the behavior of LLMs, particularly in terms of safety alignment. Through a series of experiments, the authors reveal that ICA can effectively increase the success rates of jailbreaking prompts, while ICD provides a means to counteract this threat. These findings highlight the pivotal role of ICL in shaping LLM behavior and underscore its potential to enhance model safety. In scenarios where AI agents leverage large-scale language models,, such attack and defense strategies could be adapted to secure communication channels between agents, ensuring that harmful or unintended behaviors do not arise during interactions. Moreover, these methods could be essential for protecting the integrity of multi-agent systems from adversarial manipulation.
@@
Shayegani et al. [37] introduce novel jailbreak attacks targeting vision-language models (VLMs) that integrate aligned Large Language Models (LLMs) and are resistant to text-only jailbreak techniques. The authors propose cross-modality attacks that manipulate visual and textual inputs to break the alignment of the language model. These attacks combine adversarially generated images, targeting toxic embeddings within the vision encoder, with generic textual prompts to exploit the VLM’s alignment. The images are designed using a new embedding-space-based methodology, which does not require direct access to the LLM model but only to the vision encoder. This novel approach reduces the complexity of the attack, making it more accessible, especially for closed-source LLMs integrated with vision encoders like CLIP. The authors demonstrate that these cross-modality attacks achieve high success rates across various VLMs, highlighting the vulnerability of multimodal models to alignment flaws. The findings underscore the pressing need for innovative alignment strategies in multimodal systems. For LLM-based AI agents, such attacks could significantly impact communication between agents that rely on text and visual inputs, creating new challenges in maintaining secure and reliable interactions in multimodal environments. Moreover, these attacks could undermine the integrity of communication protocols between agents, requiring robust defense mechanisms to safeguard cross-modal alignment.
@@
Anil et al. [38] explore a novel class of long-context attacks targeting large language models (LLMs), where the attack involves prompting the model with hundreds of demonstrations of undesirable behavior. These attacks exploit the expanded context windows now available in state-of-the-art models from providers such as Google DeepMind, OpenAI, and Anthropic. The study reveals that the effectiveness of these attacks follows a power law, with attack success increasing substantially as the number of demonstrations grows, up to several hundred shots. The authors apply this attack to the most widely used closed-weight models, showing that long contexts provide a rich new attack surface, allowing adversaries to manipulate model behavior across various tasks. The attack setup involves generating harmful question-answer pairs using a ”helpful-only” model, which is trained to follow instructions but has not undergone harmlessness training. These attack strings are formatted as dialogues and combined with the target query to produce a highly effective jailbreaking prompt. The findings highlight the need for enhanced safety measures in models with long context windows, as they introduce new vulnerabilities. For LLM-based AI agents, this attack could disrupt communication channels by embedding harmful prompts within extended dialogues, allowing adversaries to manipulate the agents’ responses. To safeguard such systems, novel defense mechanisms are required to handle and detect the inclusion of harmful context in multi-turn interactions.
@@
III-B3 Automated Jailbreak Prompt Generation

### v2

@@ III Input Manipulation Attacks
This section groups techniques that subvert LLM behavior by tampering with inputs — whether text, code, or multimodal data — before they reach the core model. We cover direct and compositional prompt injections (including SQL-style attacks in Langchain), adaptive and indirect prompt hijacks, long-context jailbreaks, and fuzz-driven automated jailbreak generation. Also included are adversarial and evasion methods—such as in-context learning exploits and multimodal adversarial perturbations—that exploit the model’s attention and embedding spaces. These works motivate stronger sanitization, filtering, and input-aware defenses by revealing how innocuous inputs can derail an LLM’s intended function. Tab. III presents a summary of research works on input manipulation attacks. Fig. 3 organizes input manipulation attacks into three broad categories: Prompt-Based, Jailbreaking, and Adversarial & Evasion.
III-A Formal Definition of the Input Manipulation Threat Model
To provide a rigorous foundation for the taxonomy of input manipulation threats discussed in this section, we introduce a formal mathematical definition of the Input Manipulation Threat Model. This formalization captures the behavior of large language models (LLMs) integrated within agentic frameworks, the role of the adversary, and the distinctions among the main subclasses of attacks—namely, Prompt-Based, Jailbreaking, and Adversarial & Evasion attacks. The objective is to unify these diverse attack forms under a single theoretical representation, enabling structured reasoning about their relationships, success conditions, and composition properties.
III-A1 System Model
Let $M_{\theta}$ denote a large language model (LLM) parameterized by $\theta$ and deployed within an agentic framework $\mathcal{S}=(M_{\theta},\mathcal{T},\Phi)$, where $\mathcal{T}$ represents the set of callable tools and $\Phi$ the orchestration logic (planner, memory, or retrieval). At inference time, the system processes a context:
$C=(h,x,k),$ (1)
where $h$ represents prior dialogue, $x$ denotes the current user input, and $k$ corresponds to external or retrieved knowledge. The system produces:
$(y,a)=F_{\mathcal{S}}(C),$ (2)
where $y$ is the generated output and $a\in\mathcal{A}^{*}$ represents any invoked tool actions. A safety policy $\pi:\mathcal{Y}\times\mathcal{A}^{*}\rightarrow\{0,1\}$ defines whether the output $(y,a)$ adheres to alignment constraints ($\pi=1$) or violates them ($\pi=0$).
This system model formalizes how the agent processes information and interacts with external tools. By explicitly defining the context $C=(h,x,k)$, we capture the full decision surface that an adversary may attempt to manipulate. The inclusion of tool actions $a$ is critical in agentic LLMs, as the attack surface extends beyond textual outputs to actions with real-world impact. The safety policy $\pi$ provides a binary abstraction for alignment constraints, enabling unified treatment of guardrails, refusal conditions, and tool-use restrictions.
III-A2 Adversary Model
The adversary perturbs the input context through a transformation:
$g:(h,x,k)\mapsto(h^{\prime},x^{\prime},k^{\prime}),$ (3)
with the goal of subverting the model’s intended behavior. An attack is successful if the resulting output violates the safety policy or achieves a specific malicious objective $o$, defined by:
$\pi(y,a)=0\quad\text{or}\quad\mathcal{L}(y,o)\leq\varepsilon,$ (4)
where $(y,a)=F_{\mathcal{S}}(g(h,x,k))$ and $\mathcal{L}$ represents an objective-specific loss function.
This formalization shows that the adversary may modify past conversation, the current input, or retrieved knowledge—capturing a wide range of real-world manipulation pathways. The abstraction allows the attack to be goal-driven (produce a desired output) or violation-driven (break alignment), and accommodates multi-step, indirect, and compositional adversarial strategies. This makes the model expressive enough to unify prompt injections, retrieval poisoning, and multi-stage manipulation strategies under a common framework.
III-A3 Prompt-Based Attacks
Prompt-based attacks modify natural-language or code-level instructions:
$g(h,x,k)=(h\!\oplus\!\delta_{h},\,x\!\oplus\!\delta_{x},\,k),$ (5)
where $\delta_{h}$ and $\delta_{x}$ are adversarial textual injections. The attacker aims to maximize the likelihood of unsafe or attacker-aligned outcomes:
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\max_{\delta_{h},\delta_{x}}\Pr\!\left[\pi\!\left(F_{\mathcal{S}}(h\!\oplus\!\delta_{h},\,x\!\oplus\!\delta_{x},\,k)\right)=0\ \text{or}\ \mathcal{L}\!\left(y,o\right)\leq\varepsilon\right].$} (6)
This category includes direct prompt injections, Prompt-to-SQL (P2SQL) attacks, and compositional or indirect prompt injections embedded in retrieved content.
Prompt-based attacks operate at the semantic level, where the adversary reshapes the agent’s reasoning through carefully crafted natural-language perturbations. These perturbations may be explicit (direct harmful commands) or subtle (contextual manipulation, multi-turn setup, or retrieval poisoning). By explicitly optimizing over $\delta_{h}$ and $\delta_{x}$, we show that the adversary seeks the most influential textual change that induces unsafe or malicious behavior. This general formulation captures many real-world prompt manipulation techniques used in jailbreak benchmarks, indirect injections, and downstream tool manipulation.
III-A4 Jailbreaking Attacks
Jailbreaking attacks constitute a specialized subset of prompt-based attacks focused on defeating alignment or refusal mechanisms:
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\max_{\delta}\Pr\!\left[F_{\mathcal{S}}(h,x\!\oplus\!\delta,k)\in\mathcal{Y}_{\text{forbidden}}\ \wedge\ \pi(y,a)=0\right].$} (7)
where $\delta$ is a natural-language perturbation. Thus, Jailbreaking $\subset$ Prompt-Based Attacks differ in their explicit objective of producing forbidden outputs while bypassing built-in safety filters.
Jailbreaking attacks explicitly target safety mechanisms, refusal layers, and alignment constraints. Unlike general prompt-based attacks, jailbreaks aim to induce outputs in a predefined forbidden set $\mathcal{Y}_{\text{forbidden}}$. This makes them an important diagnostic tool for evaluating whether an LLM’s safety constraints are robust or can be circumvented by carefully crafted language patterns. Recent benchmarks demonstrate that jailbreak attacks often exploit predictable guardrail weaknesses, underscoring the need for formal modeling.
III-A5 Adversarial Example & Evasion Attacks
Adversarial and evasion attacks modify low-level token or feature representations within a constrained perturbation budget. The correct formulation is:
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\min_{\Delta}\ d((h,x,k),(h^{\prime},x^{\prime},k^{\prime}))\quad\text{s.t.}\quad F_{\mathcal{S}}(h^{\prime},x^{\prime},k^{\prime})\neq F_{\mathcal{S}}(h,x,k),$} (8)
where $\Delta$ governs the allowed embedding-level perturbations and $d(\cdot,\cdot)$ measures similarity (e.g., edit distance or $\ell_{p}$ token norms). Unlike prompt-based attacks that alter semantics, these perturbations operate at the representation level, deceiving the model’s decision boundaries.
These attacks exploit the model’s sensitivity to small but carefully crafted perturbations in the embedding or token space. Even when such perturbations are imperceptible to humans, they can cause drastic behavioral changes in the model. This formulation highlights that adversarial examples target geometric weaknesses in the underlying learned representations, rather than semantic vulnerabilities. In agentic systems, even minor perturbations can alter tool invocation, decision-making, or action execution in significant and potentially harmful ways.
III-A6 Relationship Among Attack Classes
Prompt-based attacks represent the broad family of input manipulation threats. Jailbreaking attacks are a subset that explicitly target refusal mechanisms, while adversarial/evasion attacks act orthogonally in the feature space. Combined attacks can be modeled as sequential compositions:
$g_{\text{composite}}=g_{\text{adv}}\circ g_{\text{prompt}},$ (9)
enabling adversaries to exploit both semantic and embedding-level vulnerabilities simultaneously. Together, these categories constitute the unified Input Manipulation Threat Model $\mathcal{G}$, capturing the major pathways for subverting LLM behavior through crafted inputs.
This hierarchy shows that input manipulation attacks span a spectrum from high-level semantic perturbations to low-level embedding modifications. Prompt-based attacks directly influence reasoning, adversarial examples exploit geometric vulnerabilities, and jailbreaks target safety constraints. Sequential and composite attacks reflect real-world strategies where adversaries combine these dimensions to bypass increasingly complex defenses. This unified model therefore provides a comprehensive and principled way to reason about attack taxonomies, compositions, and defensive coverage.
III-B Prompt-Based Attacks
@@
III-B1 Direct Prompt Injection
@@
Perez et al. [perez2022ignore] investigate vulnerabilities in transformer-based large language models (LLMs), specifically GPT-3, that arise from malicious user interactions. Despite widespread adoption of such models in customer-facing applications, there is limited research on their susceptibility to adversarial manipulation. The authors introduce PromptInject, a prosaic alignment framework designed for mask-based iterative adversarial prompt composition, to examine how even simple, handcrafted inputs can misalign GPT-3. The study focuses on two primary attack types: goal hijacking, in which the attacker manipulates the model to perform unintended actions, and prompt leaking, in which sensitive information is inadvertently exposed through the model’s responses. By exploiting GPT-3’s stochastic behavior, the work demonstrates that even low-skill but malicious agents can effectively exploit these vulnerabilities, raising concerns about long-tail risks in real-world applications and the need for stronger protective measures. Fig. 4 illustrates the adversarial inputs designed to manipulate the instructions provided to a language model, which is presented in [perez2022ignore]. Both depicted attacks aim to subvert the prompt’s initial intent. In the objective-hijacking scenario, the attacker aims to have the model output a predefined malicious phrase, thereby overriding the original instructions. Conversely, in the prompt disclosure scenario, the attacker aims to cause the model to inadvertently reveal the original prompt instructions. The Original Application Instructions contain placeholders user_provided{user\_provided} that represent regular user inputs intended for correction. Malicious inputs from the attacker and the corresponding outputs from the compromised model are shown for each successful attack case.
III-B2 Prompt-to-SQL (P2SQL) injection attack
@@
Pedro et al. [pedro2023prompt] address security vulnerabilities in web applications that use Large Language Models (LLMs), specifically focusing on SQL injection attacks enabled by unsanitized user prompts. In systems that integrate LLMs with middleware such as LangChain, user input is converted into SQL queries that guide the model’s responses. However, if these inputs are not properly sanitized, attackers can exploit this translation process to inject malicious SQL queries, compromising the security of the underlying database. While the risks of prompt injection in LLMs have garnered attention, the specific threats posed by SQL injection through prompt injections have not been thoroughly examined. The authors present a comprehensive study of prompt-to-SQL (P2SQL) injection attacks in LangChain-based web applications. Through detailed case studies, the paper characterizes various forms of P2SQL attacks and evaluates their impact on application security. Additionally, the paper assesses seven state-of-the-art LLMs and finds that they are highly susceptible to P2SQL attacks. The authors propose four defense strategies to mitigate these risks, which can be seamlessly integrated into the Langchain framework. These defenses are validated through experimental evaluations, demonstrating their efficacy against P2SQL injection attacks in real-world applications.
III-B3 Indirect & Compositional Prompt Injection
@@
By manipulating images and sounds, Bagdasaryan et al. [bagdasaryan2023abusing] explore the potential for indirect prompt and instruction injection in multimodal large language models (LLMs). In this approach, an attacker creates an adversarial perturbation that aligns with a specific prompt or instruction. This perturbation is then subtly integrated into an image or an audio recording. When the user interacts with the model using the unaltered image or audio, the embedded perturbation influences the model’s output, effectively steering the conversation or responses toward the attacker’s preferences. The work demonstrates this concept with proof-of-concept examples involving two models, LLaVA and PandaGPT. This attack highlights the vulnerability of multimodal models to adversarial inputs, underscoring the need for stronger defenses to detect and mitigate such manipulations in real-world applications.
III-B4 Adaptive indirect prompt injection attack
@@
Zhan et al. [zhan2025adaptive] examine the security risks posed by indirect prompt injection (IPI) attacks in Large Language Model (LLM) agents that utilize external tools to interact with their environments. While various defenses have been proposed to mitigate IPI attacks, their effectiveness against adaptive attacks has not been adequately tested. The authors evaluate eight existing defense mechanisms and demonstrate that all can be bypassed by adaptive attack strategies, resulting in an attack success rate exceeding 50% in each case. This highlights significant vulnerabilities in current defense approaches and underscores the need to incorporate adaptive attack evaluation into the design of security measures for LLM agents. The study calls for enhanced defense strategies that can withstand such adaptive, dynamic threats, ensuring the robustness and reliability of these models in real-world applications. Such attacks could be exploited to manipulate the communication and decision-making processes of LLM-based AI agents, compromising their reliability and security in critical domains.
III-B5 Toxic Agent Flow attack
@@
The Toxic Agent Flow attack is a critical vulnerability in the widely used GitHub MCP server, discovered by Invariant Labs [milanta2025githubmcp], which allows attackers to hijack a user’s agent through malicious GitHub issues, leading to the leakage of private repository data. As shown in Fig. 5, the vulnerability exploits a Toxic Agent Flow, a scenario in which agents are manipulated into performing unintended actions, such as leaking sensitive data, through seemingly benign instructions. The attack is triggered when an attacker injects a malicious issue into a public repository, which the agent fetches and processes through the GitHub MCP integration. Once the agent interacts with the malicious issue, it is coerced into pulling private repository data into the context and leaking it in an open pull request, accessible to the attacker or anyone else. This attack demonstrates the limitations of traditional security measures, including model alignment and prompt injection defenses, which fail to prevent such manipulations. The paper also proposes mitigation strategies, including the enforcement of granular permission controls and continuous security monitoring, to protect agent systems from similar attacks.
@@ III-C Jailbreaking Attacks
The attack formulation [wei2023jailbreak] focuses on prompting the model to predict the next token in a sequence by using a probability distribution
@@
III-C1 Compositional Instruction Attack
@@
The recent study by Jiang et al. [jiang2025deceiving] on Compositional Instruction Attacks (CIA) presents a novel framework to better understand jailbreak vulnerabilities in large language models (LLMs), particularly in critical applications such as autonomous driving (AD). The researchers introduce two new methods for automatically generating tailored jailbreak prompts that bypass LLM security mechanisms by embedding harmful instructions within seemingly harmless ones. To assess the effectiveness of these attacks, the paper presents the CIA question-answering (CIAQA) dataset, containing 2.7K multiple-choice questions based on 900 successful jailbreaks. The experimental results reveal that these attacks can achieve a high attack success rate (ASR) of 95% or higher on popular LLMs, such as GPT-4, GPT-3.5, and Llama2-70b-chat. The study also identifies key reasons LLM defenses fail against such attacks, primarily the inability of LLMs to properly discern underlying harmful intent and to prioritize tasks effectively. To counter these threats, the authors propose an intent-based defense paradigm (IBD) that leverages LLMs’ ability to identify harmful intents, significantly reducing the attack success rate by over 74%. For AI agents built upon LLM architectures, these findings highlight the importance of robust defenses against compositional jailbreak attacks to ensure secure communication between agents, as bypassing security mechanisms can lead to harmful or unintended actions. As LLMs are integrated into autonomous agent communication protocols, similar attack vectors could be exploited, underscoring the need for stronger safeguards to protect against malicious input manipulation.
III-C2 Jailbreak via In-Context Demonstrations
@@
Wei et al. [wei2023jailbreak] explore the use of In-Context Learning (ICL) as a mechanism to influence the alignment of Large Language Models (LLMs), focusing on mitigating safety concerns related to harmful content generation. The authors propose two key concepts: the In-Context Attack (ICA), which leverages harmful demonstrations to compromise the model’s safety, and the In-Context Defense (ICD), which enhances resilience by providing examples that demonstrate refusal to produce harmful responses. The paper presents a theoretical framework that shows how a small set of carefully chosen in-context demonstrations can significantly alter LLM behavior, particularly in terms of safety alignment. Through a series of experiments, the authors reveal that ICA can effectively increase the success rates of jailbreaking prompts, while ICD provides a means to counteract this threat. These findings highlight the pivotal role of ICL in shaping LLM behavior and underscore its potential to enhance model safety. In scenarios where AI agents leverage large-scale language models, such attack and defense strategies could be adapted to secure communication channels between agents, ensuring that harmful or unintended behaviors do not arise during interactions. Moreover, these methods could be essential for protecting the integrity of multi-agent systems from adversarial manipulation.
@@
Shayegani et al. [shayegani2023jailbreak] introduce novel jailbreak attacks targeting vision-language models (VLMs) that integrate aligned Large Language Models (LLMs) and are resistant to text-only jailbreak techniques. The authors propose cross-modality attacks that manipulate visual and textual inputs to disrupt the language model’s alignment. These attacks combine adversarially generated images targeting toxic embeddings in the vision encoder with generic textual prompts to exploit the VLM’s alignment. The images are generated using a new embedding-space-based methodology that does not require direct access to the LLM model, only to the vision encoder. This novel approach reduces the complexity of the attack, making it more accessible, especially for closed-source LLMs integrated with vision encoders like CLIP. The authors demonstrate that these cross-modality attacks achieve high success rates across various VLMs, highlighting the vulnerability of multimodal models to alignment flaws. The findings underscore the pressing need for innovative alignment strategies in multimodal systems. For LLM-based AI agents, such attacks could significantly affect communication between agents that rely on both text and visual inputs, creating new challenges for maintaining secure, reliable interactions in multimodal environments. Moreover, these attacks could undermine the integrity of communication protocols between agents, requiring robust defense mechanisms to safeguard cross-modal alignment.
@@
Anil et al. [anil2024many] explore a novel class of long-context attacks targeting large language models (LLMs), in which the attack prompts the model with hundreds of demonstrations of undesirable behavior. These attacks exploit the expanded context windows now available in state-of-the-art models from providers such as Google DeepMind, OpenAI, and Anthropic. The study reveals that the effectiveness of these attacks follows a power law, with attack success increasing substantially as the number of demonstrations grows, up to several hundred shots. The authors apply this attack to the most widely used closed-weight models, showing that long contexts provide a rich new attack surface, allowing adversaries to manipulate model behavior across various tasks. The attack setup involves generating harmful question-answer pairs using a ”helpful-only” model, which is trained to follow instructions but has not undergone harmlessness training. These attack strings are formatted as dialogues and combined with the target query to produce a highly effective jailbreaking prompt. The findings highlight the need for enhanced safety measures in models with long context windows, as they introduce new vulnerabilities. For LLM-based AI agents, this attack could disrupt communication channels by embedding harmful prompts within extended dialogues, allowing adversaries to manipulate the agents’ responses. To safeguard such systems, novel defense mechanisms are needed to detect and handle the inclusion of harmful context in multi-turn interactions.
@@
III-C3 Automated Jailbreak Prompt Generation